# Thermal cycling - evidence for a generalized tunneling model and a tool to distinguish noise sources in quantum circuits

### v1

@@ Abstract
The experimental and theoretical verification of the generic existence of two types of TLSs in amorphous solids, differing by the magnitude of their interaction with the strain, and possibly by the magnitude of their dipole moment, would be a significant advance of our understanding of the nature of TLSs in amorphous solids; and with that the characteristics of noise, relaxation, decoherence, and loss in quantum circuits, and the essential physics of the low temperature universality in amorphous solids. Detection of the two types of TLSs through the dielectric loss of resonators under electric bias sweep [36] requires the strongly interacting TLSs to have also a large electric dipole moment. Detection of strongly interacting TLSs not having large electric dipoles in similar bias sweep experiments requires coupling to their elastic moment, by measuring internal friction and using an acoustic bias sweep field. Here we suggest an alternative for the detection of strongly interacting TLSs in generic amorphous solids, through thermal cycling experiments. While initialization of the TLS bath by thermal cycling is a routine procedure, careful data collecting of TLS statistics as reported by Shalibo et. al. [39], performed over variety of materials, and compared against theory, can further probe the general existence of two types of TLSs in amorphous solids, differing essentially by their interaction with the strain, as suggested by the Two-TLS model.
@@
I RATIONALE AND JUSTIFICATIONS FOR THE EXTRAPOLATION METHODOLOGY
@@
In the main part, we conducted extrapolation on the system size to assess the overall impact of the heating and cooling cycles on the excitation energies of the $\tau$ TLSs. Within the extrapolation process we utilize the probability of flip for the $S$-TLSs, $P_{s}(n)$, obtained from simulations performed on a small-scale system. In this procedure we assume that the sizes used for the direct simulation, $L_{i}=10,12,16$, suffice to produce relevant distributions $P_{s}(n)$ for extrapolation. This approximation is supported by the results presented in Fig. 1 (b) at the main part, yet we would like to look into this approximation in more detail. While TLS-TLS interaction is long range, its value at distances larger than $L/2=6$ is roughly $5$K. Since $S$-TLSs have excitation energies that are typically much larger than $5$K, and are pseudo-gapped at low energies ${\lesssim}10$K we expect that upon thermal cycling to temperature ${\lesssim}20$K the correlation of $S$-TLS flips of distances larger than $L/2=6$ to be very small.
@@
Figure S1: (a) Comparison between the distance distribution of TLS pairs in a $12^{3}$ cube with a particle density of $\rho=1/2$, and the distribution of the distances between the $S$ TLSs pairs which flipped, in the case where only $2S$ TLSs flipped. (b) Initial energies of $\tau$-TLSs that flipped after thermal cycling to different heating temperatures, for simulations where non of the $S$ TLSs have flipped.
@@
Support to this assertion is given in Fig. S1(a), where we plot for $L_{i}=12$ the distribution of distances between flipped $S$-TLS pairs in those samples where exactly two $S$-TLSs are flipped following thermal cycling. Results clearly indicates that the effect of correlated flips of $S$-TLSs is a local phenomenon.
@@
Another point to consider is that in the extrapolation process, we solely consider the number of $S$-TLSs that flipped in each cube due to the heating and cooling cycle, while disregarding the contribution of the flipped $\tau$-TLSs to the change of the excitation energies of other $\tau$-TLSs. The justification for this neglect can be divided into two scenarios. First, in cubes where at least a single $S$-TLS has flipped, the relative contribution to the change of $\tau$-TLS excitation energies by flipped $\tau$-TLSs is negligible, as $J_{0}^{\tau\tau}\ll J_{0}^{S\tau}$.
Then, in cubes where no $S$-TLSs have flipped as a result of the heating and cooling cycle, flipped $\tau$-TLSs are predominantly thermal TLSs, causing small jitter in the excitation energies of other $\tau$-TLSs, yet no long-time change in the mean of their excitation energies. In Fig. S1(b) we plot the integrated number of flipped $\tau$-TLSs as function of their excitation energies, indeed confirming that they are predominantly thermal.
II CONVERGENCE OF THE EXTRAPOLATION RESULTS
@@
In the main section, we presented the key results regarding the number of $\tau$ TLSs that experienced a change in their excitation energy beyond a specific value referred to as $\Delta E$. We also provided a detailed explanation of the extrapolation process. Fig S2 illustrates the results of the extrapolation for increasing sample sizes until convergence is reached.
@@
Figure S2: Portion of $\tau$ TLSs whose excitation energy has changed by more than $\Delta E$ = 0.1K. $L=12$ from direct calculation. Results for larger sizes are based on the probability distribution $P_{s}$ generated from the simulation for $L=12$.
@@
Figure S3: The fraction of $\tau$ TLSs whose excitation energies change by more than $\Delta E=0.1\text{K}$ for (a) different cutoff values, (b) different strain interaction ratios, and (c) different $\rho$ (TLS densities). Presented results are after extrapolation to large sizes until conversion (as in Fig. 1(b) in the main text), using $P_{s}$ deduced from simulations of size $L=12$ ($L=15$ for $\rho=0.25$, to maintain the same number of particles).
III SENSITIVITY TO PARAMETERS CHANGE
@@
Our simulations for the Two-TLS model involve several parameters: $\tilde{a}$, the cutoff distance, $g$, the ratio between the strain coupling strengths of $\tau$-TLSs and that of $S$-TLSs, and the density $\rho$ at which we embed the TLSs in the lattice. Here, we evaluate the simulation results for variations of these parameters. In all cases examined, the density of states (DOS) remains Gaussian [1], with only its width changing according to:
@@
To maintain a constant Gaussian width, we adjusted the value of the parameter $J_{0}$ accordingly.
Results, after extrapolation to large sizes until convergence, using $P_{s}$ generated from simulations of $L=12$ lattice sizes ($L=15$ in the case of $\rho=0.25$, to maintain the same number of particles in the sample), are presented in Fig. S3. We find that lowering $\tilde{a}$, lowering $g$, and lowering $\rho$, all have the effect of lowering the number of $\tau$-TlSs that change their energy by more than $0.1$K upon thermal cycling. Yet, all these changes are quantitative, leaving our qualitative result intact, i.e. at cycling temperatures $\approx 20$K the excitation energies of almost all $\tau$-TLSs change by more than $0.1$K, while the same is true for only a small number of $\tau$-TLSs at cycling temperatures $\approx 1$K.

### v2

@@ Abstract
The experimental and theoretical verification of the generic existence of two types of TLSs in amorphous solids, differing by the magnitude of their interaction with the strain, and possibly by the magnitude of their dipole moment, would be a significant advance of our understanding of the nature of TLSs in amorphous solids; and with that the characteristics of noise, relaxation, decoherence, and loss in quantum circuits, and the essential physics of the low temperature universality in amorphous solids. Detection of the two types of TLSs through the dielectric loss of resonators under electric bias sweep [36] requires the strongly interacting TLSs to have also a large electric dipole moment. Detection of strongly interacting TLSs not having large electric dipoles in similar bias sweep experiments requires coupling to their elastic moment, by measuring internal friction and using an acoustic bias sweep field. Here we suggest an alternative for the detection of strongly interacting TLSs in generic amorphous solids, through thermal cycling experiments. While initialization of the TLS bath by thermal cycling is a routine procedure, careful data collecting of TLS statistics as reported by Shalibo et. al. [39], performed over variety of materials, and compared against theory, can further probe the general existence of two types of TLSs in amorphous solids, differing essentially by their interaction with the strain, as suggested by the Two-TLS model. Furthermore, taking advantage of the quantitative dependence of our results on the parameters of the Two-TLS model (see details in Supplemental Material[42]), comprehensive study of the statistics of the changes of TLS excitation energies following thermal cycling to different temperatures would allow the extraction of the parameters of the Two-TLS model.
@@
I The Two-TLS model in disordered lattices and in amorphous solids
The low-temperature universal phenomena in amorphous solids is in fact a broad phenomena in strongly disordered systems, observed also in disordered crystals, polymers, and quasicrystals [3]. For example, KBr1-x:CNx, perhaps the archetypal disordered lattice exhibiting the universal phenomena, has, for $0.70.25<x<$ 0.7, and below $1$K, specific heat $\propto T^{\alpha}$ with $\alpha\sim 1$, thermal conductivity $\propto T^{\beta}$ with $\beta\sim 2$, and nearly temperature independent internal friction. The observation of the low-temperature universal phenomena in a large variety of disordered lattices led to the idea that the phenomena are not related to the lack of long range order in amorphous solids, but rather to the strong local disorder, shared by both amorphous solids and disordered lattices. Additional evidence for this notion was given by experiment on ion-implanted crystalline silicon. Implantation by Boron ions at high dosage leads to amorphicity, while implantation by Silicon ions keeps long-range lattice order intact. Yet, both samples show at high implantation dose the same universal phenomena, suggesting that local disorder rather than the absence of long-range order dictate the low-temperature phenomena in strongly disordered and amorphous solids. This led the authors to argue that the defects in the crystal should be used to model the excitations in the amorphous silicon, rather than the amorphous structure itself [44]. Disordered lattices are advantageous in the study of the low-temperature universal phenomena for various reasons. TLSs in disordered solids can be identified, e.g. CN dipoles in KBr1-x:CNx, and disordered lattices can be carefully manipulated to address physical questions. A prominent example is the debate of whether the universal phenomena is dictated by interactions between the tunneling defects, or by the strain disorder. To address this question a series of experiments studied highly strained disordered lattice, e.g. KBr0.5:Cl0.5, with a small concentration of tunneling defects, e.g. CN impurities [43, 45, 46], where it was shown that universal phenomena, linear in the concentration of the tunneling defects, appear, despite the smallness of interactions between the dilute defects, pointing to the strong strain disorder as dictating the universal low-temperature phenomena. Such experiments led Pohl et.al., [47] to argue that the nature of the defects, and the cause for their saturation density, should be explored in disordered crystals, rather than in amorphous solids.
Following this proposal, the Two-TLS model was suggested in an attempt to describe properly the low-energy properties of disordered lattices in a model which will then allow making predictions to be tested also in amorphous solids, to support its general validity in describing the low-temperature properties in strongly disordered and amorphous solids. Considering for example KBr1-x:CNx, natural candidates for tunneling states are the CN defects, which, however, can tunnel either by $90^{\circ}$ rotation or by $180^{\circ}$ flip. As a result of inversion symmetry, CN rotations and flips, however, have a very different strain interaction. In particular a single CN impurity in an otherwise pure lattice has, to first order, zero interaction of the flip TLS with the strain. In a strongly disordered KBr1-x:CNx system, with $0.70.25<x<$ 0.7, inversion symmetry is not purely maintained. Yet, the interactions of CN flips with the strain are proportional to the deviations from local inversion symmetry, which are, at near neighbor distance, of the order of $3\%$ in strongly disordred lattices, as well as in amorphous solids [48]. As a result, CN flips and rotations constitute two sets of tunneling TLSs, differing by their interaction with the strain, denoted $\gamma_{w}$ and $\gamma_{s}$ respectively, with $g\equiv\gamma_{w}/\gamma_{s}\approx 0.03$. Thus, the Two-TLS model generalizes the interaction of TLSs with the strain given by the STM
$H_{\rm TLS-ph}\;=\;\sum_{j}\sum_{\alpha,\beta}\;\gamma^{\alpha\beta}\tau_{j}^{z}\;\frac{\partial u_{j}^{\alpha}}{\partial{\bf r}_{j}^{\beta}}\,.$ (S1)
to include the two sets of TLSs differing by their strain interaction, reading:
$H^{\rm tot}_{\rm TLS-ph}\;=\;\sum_{j,\alpha,\beta}\;\gamma_{\rm s}^{\alpha\beta}S_{j}^{z}\;\frac{\partial u_{j}^{\alpha}}{\partial{\bf r}_{j}^{\beta}}+\sum_{j^{\prime},\alpha,\beta}\;\gamma_{\rm w}^{\alpha\beta}\tau_{j^{\prime}}^{z}\;\frac{\partial u_{j^{\prime}}^{\alpha}}{\partial{\bf r}_{j^{\prime}}^{\beta}}.$ (S2)
Here $u_{j}^{\alpha}$ denotes the $\alpha=x,y,z$ component of the phononic amplitude at ${\bf r}_{j}$. This bi-modality of the strain interaction was confirmed numerically using ab-initio and DFT calculations in [26]. For the STM form of the TLS-phonon interaction, Eq. (S1), integration out of the phonons results in an effective Hamiltonian for TLS-TLS interactions, given by
$\mathcal{H}_{\tau int}=-\sum_{i\neq j}\frac{1}{2}J_{ij}^{\tau\tau}\tau_{i}\tau_{j}\,.$ (S3)
and including the static disorder induced random field one obtains Eq. (3) in the main part. For the generalized Two-TLS interaction model, Eq. (S2), similar integration out of the phonons results in the effective TLS-TLS interactions given in Eq. (1) in the main part, and repeated here for convenience:
$\mathcal{H}_{s\tau}=-\sum_{i\neq j}[\frac{1}{2}J_{ij}^{ss}S_{i}S_{j}+J_{ij}^{s\tau}S_{i}\tau_{j}+\frac{1}{2}J_{ij}^{\tau\tau}\tau_{i}\tau_{j}].$ (S4)
The form of the above interaction Hamiltonian has essential consequences on the low energy DOS of the $\tau$-TLSs and of the $S$-TLSs. For the $\tau$-TLSs, the dominate energy is given by its interaction with the $S$-TLSs, which are mostly frozen at low energies, leading to a similar effective Hamiltonian to that of the STM [Eq. (3) in the main part]. However, the $S$-TLSs are pseudo-gapped at low energies, consequence of their interactions with other $S$-TLSs, and with the more abundant at low energies $\tau$-TLSs, resulting in power-low energy dependent $S$-TLS DOS at low energies [41]. This form of the DOS was calculated analytically for the model Hamiltonian in Eq.(S4) [41], as well as from first principles for the KBr1-x:CNx system, using hybrid molecular statics and Monte-Carlo numerical technique [37], with excellent agreement between the results for the real system and for the model.
As a result of the scarceness of the $S$-TLSs at low energies, their effect on most equilibrium properties is negligible. The $\tau$-TLSs are equivalent to the standard TLSs in the STM. Yet, the smallness of their interaction with the strain dictates the effective disorder exerted on the $\tau$-TLSs to be $gJ_{0}\approx 10$K, i.e. about $g$ times smaller than assumed by the STM. Even more importantly, their DOS at low energies is proportional to $(\gamma_{w}\gamma_{s})^{-1}$, leading to smallness of $\gamma_{w}/\gamma_{s}$ of the tunneling strength, which together with the standard logarithmic factor of the ratio of the maximal to minimal TLS tunneling amplitudes $\ln(\Delta_{0}^{max}/\Delta_{0}^{min})$, produces the smallness and universality of the tunneling strength. Thus, within the Two-TLS model, the universality of the low-temperature properties of strongly disordered and amorphous solids lies in the rather universal value of the typical strain in such systems, which dictates the value of $g\approx 0.03$.
Unlike the case in disordered lattices, the identification of the tunneling states in amorphous solids is a formidable task. Thus, to address the validity of the Two-TLS model for amorphous solids one resorts to making predictions unique to the Two-TLS model, and checking them against experimental results. For the Two-TLS model to give predictions which are different from that of the STM one needs to have properties with a dominant contribution of the $S$-TLSs, despite their scarceness at low energies in equilibrium. Examples are the fast switching of thermal $S$-TLSs dictating echo dephasing of resonant TLSs [34], loss of resonators under fast bias sweeping which brings $S$-TLSs from high energies into resonance [36, 38] and thermal cycling experiments, as discussed in the current manuscript.
II RATIONALE AND JUSTIFICATIONS FOR THE EXTRAPOLATION METHODOLOGY
@@
In the main part we conduct extrapolation on the system size to assess the overall impact of the heating and cooling cycles on the excitation energies of the $\tau$-TLSs. Within the extrapolation process we utilize the probability of flip for the $S$-TLSs, $P_{s}(n)$, obtained from simulations performed on a small-scale system. In this procedure we assume that the sizes used for the direct simulation, $L_{i}=10,12,16$, suffice to produce relevant distributions $P_{s}(n)$ for extrapolation. This approximation is supported by the results presented in Fig. 1 (b) in the main part, yet we would like to look into this approximation in more detail. While TLS-TLS interaction is long range, its value at distances larger than $L/2=6$ is roughly $5$K. Since $S$-TLSs have excitation energies that are typically much larger than $5$K, and are pseudo-gapped at low energies ${\lesssim}10$K we expect that upon thermal cycling to temperature ${\lesssim}20$K the correlation of $S$-TLS flips of distances larger than $L/2=6$ to be very small.
@@
Figure S1: (a) Comparison between the distance distribution of TLS pairs in a $12^{3}$ cube with a particle density of $\rho=1/2$, and the distribution of the distances between the S-TLSs pairs which flipped, in the case where only 2 S-TLSs flipped. (b) Cumulative distribution function (CDF) of the initial energies of $\tau$-TLSs that flipped after thermal cycling to different heating temperatures, in simulations where none of the S-TLSs flipped.
@@
Support to this assertion is given in Fig. S1(a), where we plot for $L_{i}=12$ the distribution of distances between flipped $S$-TLS pairs in those samples where exactly two $S$-TLSs are flipped following thermal cycling. Results clearly indicate that the effect of correlated flips of $S$-TLSs is a local phenomenon.
@@
Another point to consider is that in the extrapolation process, we solely consider the number of $S$-TLSs that flipped in each cube due to the heating and cooling cycle, while disregarding the contribution of the flipped $\tau$-TLSs to the change of the excitation energies of other $\tau$-TLSs. The justification for this neglect can be divided into two scenarios. First, in cubes where at least a single $S$-TLS has flipped, the relative contribution to the change of $\tau$-TLS excitation energies by flipped $\tau$-TLSs is negligible, as $J_{0}^{\tau\tau}\ll J_{0}^{S\tau}$. Then, in cubes where no $S$-TLSs have flipped as a result of the heating and cooling cycle, flipped $\tau$-TLSs are predominantly thermal TLSs, causing small jitter in the excitation energies of other $\tau$-TLSs, yet no long-time change in the mean of their excitation energies. In Fig. S1(b) we plot the integrated number of flipped $\tau$-TLSs as function of their excitation energies, indeed confirming that they are predominantly thermal.
III CONVERGENCE OF THE EXTRAPOLATION RESULTS
@@
In the main part we present the key results regarding the number of $\tau$-TLSs that experienced a change in their excitation energy beyond a specific value referred to as $\Delta E$. We also provide a detailed explanation of the extrapolation process. Fig S2 illustrates the results of the extrapolation for increasing sample sizes until convergence is reached.
@@
Figure S2: Fraction of $\tau$-TLSs whose excitation energy has changed by more than $\Delta E$ = 0.1K. $L=12$ from direct calculation. Results for larger sizes are based on the probability distribution $P_{s}$ generated from the simulation for $L=12$.
@@
Figure S3: Fraction of $\tau$-TLSs whose excitation energies change by more than $\Delta E=0.1\text{K}$ for (a) different cutoff values, (b) different strain interaction ratios, and (c) different $\rho$ (TLS densities). Presented results are after extrapolation to large sizes until conversion (as in Fig. 1(b) in the main part), using $P_{s}$ deduced from simulations of size $L=12$ ($L=15$ for $\rho=0.25$, to maintain the same number of particles).
IV SENSITIVITY TO PARAMETERS CHANGE
@@
Our simulations for the Two-TLS model involve several parameters: $\tilde{a}$, the interaction cutoff distance, $g$, the ratio between the strain coupling strengths of $\tau$-TLSs and that of $S$-TLSs, and the density $\rho$ at which we embed the TLSs in the lattice. Here, we compare the simulation results for variations of these parameters while keeping the $\tau$-TLS DOS unvaried. The latter is Gaussian [41], with width given by
@@
which we keep constant by adjusting the parameter $J_{0}$.
Results are presented in Fig. S3 following extrapolation to large system sizes until convergence. The calculations use $P_{s}$ generated from simulations of $L=12$ lattice sizes ($L=15$ for $\rho=0.25$, to maintain the same number of particles in the sample). The main conclusion drawn from our results here is that for parameter changes which are within the regime compatible with the Two-TLS model our qualitative results remain intact: at cycling temperatures of $\approx 20$ K, the excitation energies of almost all $\tau$-TLSs change by more than 0.1 K, whereas at cycling temperatures of $\approx 1$K, only a small number of $\tau$-TLSs exhibit such a change. Yet, the quantitative changes as function of the relevant parameters may turn out useful in extracting the model parameters from experiments.
Let us now discuss the quantitative dependence on simulation parameters of the number of $\tau$-TLSs that upon thermal cycling change their excitation energy by more than $0.1$K. While we keep the $\tau$-TLS DOS unchanged, the change of parameters has two main effects on the system: (i) changing the interaction strength between the TLSs (ii) changing the DOS of the $S$-TLSs. With regard to the former, the change of $S-\tau$ interaction strength changes the radius of influence of a flipped $S$ TLS, given by [41]
$R$ =$3]{\frac{4J_{0}g}{\Delta E}+\tilde{a}^{3}}\,.$ (S6)
For $\Delta E\ll 4J_{0}g$, we obtain
$N$ $\approx\frac{16\pi J_{0}g}{3\Delta E}\rho\,.$ (S7)
For a specific $\Delta E$, this simplifies to
$N$ $\propto J_{0}g\rho$ (S8)
and substituting $J_{0}$ from equation S5, we get
$N$ $\propto\rho^{1/2}\tilde{a}^{3/2}\,.$ (S9)
As a result, the number of $\tau$-TLSs that change their excitation energy at a given $\Delta E$ given a single flipped $S$-TLS increases with enhanced $\tilde{a}$ and with enhanced $\rho$. A competing effect is the reduction of the low energy DOS of $S$-TLSs with enhanced $\tilde{a}$ and $\rho$. Yet, this effect is subdominant. For example, the change of $\tilde{a}$ from $2$ to $1$ results in an enhancement of the $S$-TLS DOS by only $\approx 50\%$ (see Fig. $4$ in Ref. [41]). Consequently, the overall dependence of the number of $\tau$-TLSs which experience a given change in their excitation energy upon thermal cycling is monotonically increasing as function of $\tilde{a}$ and as function of $\rho$. The effect becomes less dominant with the increase of the cycling temperature, as regions within the radius of influence of distinct flipped $S$-TLSs begin to overlap.
Changing the value of $g$, however, while maintaining $E_{\text{typ}}^{\tau}$ at a constant value by keeping $J_{0}\propto 1/g$ results in a significant effect on the DOS of the $S$-TLSs. The reduction of $g$ and the corresponding increase of $J_{0}$ increases the typical value of $S$-TLS energies, and appreciably decreases the DOS of the $S$-TLSs at low energies. At the same time, from equation (S9) we find that the number of $\tau$-TLSs affected by a single $S$-TLS flip does not depend on $g$. We thus find a decrease in the total number of affected $\tau$-TLSs due to a reduction in $g$.